Randy Johnson throws a baseball horizontally from the top of a building as fast as his 2004 record of 102.0 mph (45.6 m/s). How much time passes until it moves at an angle 13.0 degrees below the horizontal? Ignore air resistance.
1
Expert's answer
2011-10-11T13:03:42-0400
Let Vx = 45.6 m/s be the horizontal component of the velocity and Vy - vertical
component. Vx doesn't change with time because we ignore air resistance.
Vy
changes with acceleration g=9.8 m/s^2. So, Vy(t) = gt. We need Vy / Vx =
tan(13.0). So, gt = Vx*tan(13.0).
t = Vx*tan(13.0)/g = 45.6*tan(13.0)/9.8 =
1.07 seconds.